# Hunting axion dark matter signatures in low-frequency terrestrial magnetic fields

### v1

@@ Abstract
We show that Earth’s natural environment can serve as a powerful probe for ultralight axion dark matter. In the presence of global geomagnetic fields, the axions with masses ranging from $10^{-15}$ eV–$10^{-13}$ eV induce electromagnetic waves in the (sub-) extremely low-frequency band ($0.3-30$ Hz) through the axion-photon coupling. We predict the amplitude of induced magnetic fields in the Earth-ionosphere cavity, taking the finite conductivity of the atmosphere into account. This allows us to constrain the axion-photon coupling parameter, $g_{\rm a\gamma}$, from the long-term monitoring data of the low-frequency magnetic fields, resulting in a significant improvement from the previous constraints down to $g_{\rm a\gamma}\lesssim 4\times 10^{-13}\,{\rm GeV}^{-1}$ for axion mass $\sim 3\times 10^{-14}\,\rm{eV}$.
preprint: YITP-25-54
@@
In this Letter, we consider a novel search for axion DM, utilizing the terrestrial EM fields in the extremely low-frequency (ELF) band. Coupled with the static geomagnetic field, coherently oscillating axions permeating the entire Earth can generate monochromatic EM waves at a frequency corresponding to the axion mass. In particular, EM waves generated near the Earth’s surface at frequencies of $1\,{\rm Hz}\lesssim f\lesssim 100$ Hz are confined between the Earth’s surface and the ionosphere, making the natural environment of the Earth in the ELF range a powerful window for exploring ultralight axions (see Ref. Beadle et al. (2024) for a possible window at the MHz band). The use of such an environment has been pointed out in Arza et al. (2022). Focusing especially on the frequency of $f\lesssim 10^{-2}$ Hz, they have placed a constraint on $g_{\rm a\gamma}$ in the mass range, $2\times 10^{-18}\,{\rm eV}\lesssim m_{\rm a}\lesssim 7\times 10^{-17}\,{\rm eV}$, utilizing the magnetometer network on Earth. Recently, the constraint has been extended to heavier masses, $m_{\rm a}\lesssim 4\times 10^{-15}\,{\rm eV}$, with high-sampling measurement data Friel et al. (2024). There are other efforts constraining the coupling with direct measurements in the mass range, $10^{-15}\,{\rm eV}$$10^{-12}\,{\rm eV}$ Anastassopoulos et al. (2017); Altenmüller et al. (2024); Sulai et al. (2023); Oshima et al. (2023); Bloch and Kalia (2024).
@@
Here, our prime focus is the EM fields at the frequencies of $f=0.3-30$ Hz, whose wavelength is comparable or longer than the Earth circumference. In contrast to the frequency range of $f\lesssim 1$ Hz studied in Refs. Arza et al. (2022); Friel et al. (2024), resonant behavior of EM waves is expected, and reliable predictions requires a proper theoretical modeling that accounts for atmospheric conductivity. The present study allows us to explore an axion signature or to put a tight bound on the axion-photon coupling across a mass range of $10^{-15}\,{\rm eV}-10^{-13}\,{\rm eV}$.
@@
Axion-induced signal in terrestrial $B$-fields. In the presence of axions, the Maxwell equations are modified, and there appear effective charge and current through the interaction $\mathcal{L}_{\rm int}$. If axions comprise the DM, they have non-relativistic velocity, $v_{\rm DM}\sim 10^{-3}$, and the terms involving spatial gradient, $|\nabla a|\sim m_{\rm a}v_{\rm DM}a$, are generally suppressed, compared to those involving the time derivative $|\partial_{t}a|\sim m_{\rm a}a$, where we adopt the natural units, and set the speed of light to unity. Then, the relevant modification appears only through the effective current, given by $\boldsymbol{J}_{\rm eff}=-g_{\rm a\gamma}\,\partial_{t}(a\,\boldsymbol{B})$. Further, the axions are predominantly coherent, and exhibit an oscillatory behavior such that $a\propto e^{-i\,m_{\rm a}t}$. This implies that when coupled with a static magnetic field, the axion fields produce an alternating current with a frequency of $f_{\rm a}=m_{\rm a}/(2\pi)\simeq 2.4\,(m_{\rm a}/10^{-14}\,{\rm eV})$ Hz, which can serve as a source of monochromatic EM waves at the same frequency $f_{\rm a}$ via the modified Ampére-Maxwell law. Assuming the characteristic scale of $R$, its magnetic field amplitude $\boldsymbol{B}_{\rm a}$ is estimated roughly by equating the terms in the Ampére-Maxwell law, ×$\nabla\times\boldsymbol{B}\sim\boldsymbol{J}_{\rm eff}$. This yields $|\boldsymbol{B}_{\rm a}|\sim g_{\rm a\gamma}m_{\rm a}a_{0}\,R\,|\boldsymbol{B}_{0}|$, with $\boldsymbol{B}_{0}$ and $a_{0}$ being respectively the amplitudes of static magnetic and axion fields.
In the terrestrial environment, the geomagnetic field serves as a representative static and global magnetic field, characterized by a dipole configuration, with its strength typically of $|\boldsymbol{B}_{\rm geo}|\sim 25-65\,\mu$T Alken et al. (2021). Setting $\boldsymbol{B}_{0}$ to $\boldsymbol{B}_{\rm geo}$ and assuming that the size of magnetic field is comparable to the Earth radius, $R_{\rm E}=6371$ km, we have
$|\boldsymbol{B}_{\rm a}|$ $\sim 0.3\,{\rm pT}\left(\frac{g_{\rm a\gamma}}{10^{-10}\,{\rm GeV}^{-1}}\right)\left(\frac{\rho_{\rm DM}}{0.3\,{\rm GeV}\,{\rm cm}^{-3}}\right)^{1/2}$
×$\times\left(\frac{R}{R_{\rm E}}\right)\left(\frac{|\boldsymbol{B}_{\rm geo}|}{50\,\mu{\rm T}}\right).$ (1)
@@
Although the amplitude at Eq. (1) is more than seven orders of magnitude smaller than that of the geomagnetic fields, it is still detectable with high-precision commercial magnetometers. Further, at $f=0.3-30$ Hz, the major background noise is a random superposition of transient EM waves (e.g., Barr et al. (2000); Simões et al. (2012)), which can be discriminated from the static axion signal having a sharp spectral line by using a long-term monitoring data of terrestrial magnetic fields.
@@
Note that Eq. (3) is derived by using the wave equation for the electric field alongside the Maxwell-Faraday equation. Here, the influence of the atmospheric conductivity is incorporated into the complex refractive index $n$ through $n^{2}=1+i\,(\sigma/m_{\rm a})$, where the conductivity $\sigma$ increases monotonically with the radial coordinate $r$, from $10^{-14}$ to $10^{-3}$ S/m between the Earth’s surface and the ionosphere (e.g., Volland (1984); Kudintseva et al. (2016); Malta and Helayël-Neto (2022)).
Figure 1: Top: Radial mode function $b_{\ell}$ for the lowest six multipoles, evaluated at the Earth’s surface $r=R_{\rm E}$. The results are independent of location, and plotted against the axion mass $m_{\rm a}$ (lower) and the induced EM wave frequency $f_{\rm a}$ (upper). Bottom: Expected amplitude of the axion-induced magnetic field at Eskdalemuir observatory, normalized by the axion-photon coupling strength of $10^{-10}\,$eV-1. Following Eq. (2), this is obtained by summing the mode functions from the top panel, weighted by the harmonic coefficients of geomagnetic field and vector spherical harmonics, and evaluating at the observatory’s location.
Adopting the conductivity profile numerically tabulated in Ref. Kudintseva et al. (2016) [^1] The induced magnetic field amplitude in Fig. 1 is largely unaffected by the choice of conductivity profiles, e.g., Galuk et al. (2018); Nickolaenko et al. (2016)., we solve the radial mode equation by treating the radial dependence of $\sigma$ or $n$ as a sum of spatially constant segments. This approach allows us to analytically solve for $b_{\ell}$, while satisfying the boundary conditions at the Earth’s surface and above the ionosphere Taruya et al. (2025). The top panel of Figure 1 plots the resultant mode functions for the lowest 6 multipoles at the Earth’s surface, given as a function of $m_{\rm a}$ or the frequency $f_{a}$ of the induced EM signal. There is a prominent peak structure in each multipole, known as the Schumann resonance Schumann (1952a, b) (see also Refs. Jackson (1998); Nickolaenko and Hayakawa (2002); Simões et al. (2012); Nickolaenko and Hayakawa (2013)), which is caused by EM waves trapped between the Earth’s surface and the ionosphere, forming a resonant cavity. The peak frequencies are close to $\sqrt{\ell(\ell+1)}/(2\pi\,R_{\rm E})\simeq 7.46\sqrt{\ell(\ell+1)}$ Hz, as predicted in the idealized case of a perfectly conducting Earth-ionosphere cavity, although the actual peak positions are systematically lower than these values by 20-30% due to losses in the propagation of EM waves. As a result, the resonant peaks have a finite amplitude, approximately $b_{\ell}\sim 5$ for $\ell=1-6$, unlike the idealized case where the amplitudes exhibit a divergent behavior.
@@
The basic data analysis is summarized as follows. First, a continuous dataset spanning more than one month is extracted from the full data, which is then divided into segments of length $T_{\rm seg}=8\,{\rm hrs}$. Second, each data segment is Fourier-transformed, and its PSD is computed. The frequency bin size of each segment is $\Delta f_{\rm a}\simeq 1/T_{\rm seg}\simeq 3.5\times 10^{-5}$ Hz, roughly corresponds to the bandwidth of the axion-induced signal at $f_{\rm a}=35$ Hz, above which the signal decays rapidly as shown in Fig. 1. Axion-induced signals at $f_{\rm a}\lesssim 35\,\,\rm{Hz}$ are well within a single frequency bin. Considering the coherence time, the segment size of 8 hrs may not be an optimal choice but is chosen so as to avoid unknown-cause short-duration transient noise with huge amplitude of $\sim 10^{8}\,\rm{pT}$ that happens irregularly. Due to the transient noise, the fraction of the data segments with good quality is roughly $\sim 10$% of the total number of the data segments, $N_{\rm seg}=9909$.
@@
Constraining axion-photon coupling. Given the PSD of the magnetic fields, we stack them with the weight of the inverse of noise variance computed from the spectrum at surrounding frequencies. Then we subtract the smooth component, obtained by applying a quadratic Butterworth filter with a sampling frequency of $0.05\,\rm{Hz}$. This allows the axion-induced line signals to add constructively, while reducing other components (magnetometer noise variance and the Schumann resonance spectrum). Dividing by the sum of the inverse-of-noise-variance weight for data segments, we obtain the weighted average differential PSD, $\hat{s}(f_{\rm a})$. The procedure above still leads to numerous lines at integer-multiple frequencies, the exact origin of which is unknown. Since these are likely artificial, we remove the data surrounding these frequencies (±$\pm 1000$ bins). The resultant PSD is shown in Fig. 2.
@@
Figure 2: Weighted-average differential PSD of the magnetic fields, stacked over all eight-hour data segments with noise-based weights and subtracted the smooth spectrum, $\hat{s}(f_{\rm a})$ in blue. The standard deviation is depicted in orange. The axion signal candidates (SNR$>3$) are marked by diamonds in cyan. The inset highlights the signal candidate with the highest SNR (13.342(13.342( 13.342).
@@
The search for the axion-induced signal consists of evaluating the signal-to-noise ratio (SNR) for each frequency bin and selecting the candidates that exceed a certain threshold. The standard deviation of the weighted-average differential PSD is calculated from the surrounding ±$\pm 200$ bins, excluding the three bins around $f=f_{\rm a}$, as shown by the orange curve in Fig. 2. We regard this as noise amplitude to estimate the SNR for the axion signal at $f_{\rm a}$, and evaluate the SNR from $0.01$ Hz, below which the number of samples is insufficient, to $44$ Hz, above which the estimated spectrum becomes less reliable because it is close to the Nyquist frequency.
@@
We then identify signal candidates by requiring (i) ${\rm SNR}>2$ in all-year data, (ii) ${\rm SNR}$ in each year’s data larger than the SNR threshold (2 in our case) weighted by the sum of the inverse of noise variances in the year, ensuring persistence over the entire observation period, and (iii) that their frequencies are at least $10^{-3}\,\rm{Hz}$ away from multiples of $0.05\,\rm{Hz}$. The third condition is for excluding likely artificial line noises. We found $342$ candidates (see also Nishizawa et al. (2025) for signal candidates with different SNR thresholds). These candidates appear in a single frequency bin, consistent with the sharpness of an axion signal. There are 1 candidate with significant SNR (13.342) and 31 candidates with SNR$>5$ for the all-year data. In Fig. 2, representative candidates of SNR$>3$ are depicted by cyan-filled diamonds. Lacking enough information to veto them, we retain them as potential axion signal candidates.
@@
We note that below $1$ Hz, Ref.  Friel et al. (2024) reported three signal candidates with modest significance using SuperMAG 1-second sampling data. Although our analysis identified five candidates with SNR$>2$, none of them matched those reported in Ref. Friel et al. (2024). Given that our measurement sensitivity is comparable to theirs [^2] To be precise, our sensitivity is slightly worse than theirs for the candidate at $0.2630$Hz identified by Ref.  Friel et al. (2024)., the signal candidates identified below $1$ Hz in both analyzes can be ruled out.
@@
Apart from the signal candidates mentioned above, no axion-like signal was found. We can therefore place upper limits on the axion coupling strength in the frequency range of $0.43-43.9$ Hz. The 95% C. L. upper limit on the axion-photon coupling strength is obtained from
@@
The $\hat{s}_{\rm obs}$ is the observed value of the weighted-average differential PSD, which can be converted to the axion coupling strength by equating it with $\langle\hat{s}_{\rm a}(m_{\rm a})\rangle\equiv 2T_{\rm seg}|\boldsymbol{B}_{\rm a}(m_{\rm a})|^{2}$ (see Ref. Nishizawa et al. (2025) for the derivation) and using the theoretical curve in the bottom panel of Fig. 1. The $p[\hat{s}(m_{\rm a})|g_{\rm a\gamma}]$ is the probability density distribution of the weighted-average differential PSD in the presence of axions, constructed from data $\hat{s}$ of the surrounding ±$\pm 200$ bins (excluding the three bins around $f=f_{\rm a}$) and shifted by the axion signal $\hat{s}_{\rm a}$ at $f=f_{\rm a}$.
@@
Figure 3 shows the upper limit on the axion coupling we obtained (in blue), along with other observation bounds. Tightest limit is obtained around an axion mass of $3\times 10^{-14}\,\rm{eV}$, which is improved by about two orders of magnitude compared to the CAST result  Anastassopoulos et al. (2017); Altenmüller et al. (2024), and exceeds the limit from astrophysical X-ray observations by Chandra Reynés et al. (2021); Reynolds et al. (2020); Dessert et al. (2022). In the lower mass range, our constraint is weakened because of fixing the data segment size to 8 hrs and using broader frequency resolution, irrespective of the coherence time of axions. In the higher mass range above $10^{-14}\,\rm{eV}$, the constraint is also weakened due to the suppression of the magnetic field’s response to axions.
@@
Figure 3: Constraint (95% C. L.) on the axion-photon coupling $g_{\rm a\gamma}$ from the long-term monitoring data of magnetic fields at Eskdalemuir observatory (this work, blue) except for the notched frequency bands. Other excluded regions are from SuperMAG Arza et al. (2022); Friel et al. (2024), SNIPE Sulai et al. (2023), Planck and unWISE blue galaxy sample Goldstein et al. (2024), CAST Anastassopoulos et al. (2017); Altenmüller et al. (2024), and Chandra Reynés et al. (2021); Reynolds et al. (2020); Dessert et al. (2022). The data are taken from Axi .
@@
Conclusions and outlook. In this Letter, we properly accounted for atmospheric conductivity to predict terrestrial EM waves induced by coherently oscillating axions, enabling a search for axion DM signals in public magnetic field data. We set the most stringent direct upper bound on the axion-photon coupling parameter at $10^{-15}\,{\rm eV}\lesssim m_{\rm a}\lesssim 10^{-13}\,{\rm eV}$, and identify potential signal candidates with significant SNR for further study. Follow-up ground-based experiments like DANCE Obata et al. (2018) or X-ray observations by Athena Sisk-Reynés et al. (2023) will be crucial to verify or rule out these candidates. While we focus on axions as a representative ultralight DM, our methodology applies to other DM coupled to EM fields, such as dark photon DM Nelson and Scholtz (2011); Fabbrichesi et al. (2020), whose kinetic mixing can be constrained using the same dataset.

### v2

@@ Abstract
We show that Earth’s natural environment can serve as a powerful probe for ultralight axion dark matter. In the presence of global geomagnetic fields, the axions with masses ranging from $10^{-15}$ $10^{-13} eV$ induce electromagnetic waves in the (sub-) extremely low-frequency band (0.3300.3-30 Hz) through the axion-photon coupling. We predict the amplitude of induced magnetic fields in the Earth-ionosphere cavity, taking the finite conductivity of the atmosphere into account. This allows us to constrain the axion-photon coupling parameter, $g_{\rm a\gamma}$, from the long-term monitoring data of the low-frequency magnetic fields, resulting in a significant improvement from the previous constraints down to $g_{\rm a\gamma}\lesssim 4\times 10^{-13}\,{\rm GeV}^{-1}$ for axion mass $\sim 3\times 10^{-14}\,\rm{eV}$.
[…]
xxxx, xxx
@@
In this Letter, we consider a novel search for axion DM, utilizing the terrestrial EM fields in the extremely low-frequency (ELF) band. Coupled with the static geomagnetic field, coherently oscillating axions permeating the entire Earth can generate monochromatic EM waves at a frequency corresponding to the axion mass. In particular, EM waves generated near the Earth’s surface at frequencies of $1\,{\rm Hz}\lesssim f\lesssim 100$ Hz are confined between the Earth’s surface and the lower ionosphere, making the natural environment of the Earth in the ELF range a powerful window for exploring ultralight axions (see Ref. Beadle et al. (2024) for a possible window at the MHz band). The use of such an environment has been pointed out in Arza et al. (2022). Focusing especially on the frequency of $f\lesssim 10^{-2}$ Hz, they have placed a constraint on $g_{\rm a\gamma}$ in the mass range, $2\times 10^{-18}\,{\rm eV}\lesssim m_{\rm a}\lesssim 7\times 10^{-17}\,{\rm eV}$, utilizing the magnetometer network on Earth. Recently, the constraint has been extended to heavier masses, $m_{\rm a}\lesssim 4\times 10^{-15}\,{\rm eV}$, with high-sampling measurement data Friel et al. (2024). There are other efforts constraining the coupling with direct measurements in the mass range, $10^{-15}\,{\rm eV}$$10^{-12}\,{\rm eV}$ Anastassopoulos et al. (2017); Altenmüller et al. (2024); Sulai et al. (2023); Oshima et al. (2023); Bloch and Kalia (2024).
@@
Here, our prime focus is the EM fields at the frequencies of f=0.330f=0.3-30 Hz, whose wavelength is comparable or longer than the Earth circumference. In contrast to the frequency range of $f\lesssim 1$ Hz studied in Refs. Arza et al. (2022); Friel et al. (2024), resonant behavior of EM waves is expected, and reliable predictions requires a proper theoretical modeling that accounts for atmospheric conductivity. The present study allows us to explore an axion signature or to put a tight bound on the axion-photon coupling across a mass range of $10^{-15}\,{\rm eV}-10^{-13}\,{\rm eV}$.
@@
Axion-induced signal in terrestrial $B$-fields. In the presence of axions, the Maxwell equations are modified, and there appear effective charge and current through the interaction $\mathcal{L}_{\rm int}$. If axions comprise the DM, they have non-relativistic velocity, $v_{\rm DM}\sim 10^{-3}$, and the terms involving spatial gradient, $|\nabla a|\sim m_{\rm a}v_{\rm DM}a$, are generally suppressed, compared to those involving the time derivative $|\partial_{t}a|\sim m_{\rm a}a$, where we adopt the natural units, and set the speed of light to unity. Then, the relevant modification appears only through the effective current, given by $\boldsymbol{J}_{\rm eff}=-g_{\rm a\gamma}\,(\partial_{t}a)\,\boldsymbol{B}$. Further, the axions are predominantly coherent, and exhibit an oscillatory behavior such that $a\propto e^{-i\,m_{\rm a}t}$. This implies that when coupled with a static magnetic field, the axion fields produce an alternating current with a frequency of $f_{\rm a}=m_{\rm a}/(2\pi)\simeq 2.4\,(m_{\rm a}/10^{-14}\,{\rm eV})$ Hz, which can serve as a source of monochromatic EM waves at the same frequency $f_{\rm a}$ via the modified Ampére-Maxwell law. Assuming the characteristic scale of $R$, its magnetic field amplitude $\boldsymbol{B}_{\rm a}$ is estimated roughly by equating the terms in the Ampére-Maxwell law, $\nabla\times\boldsymbol{B}\sim\boldsymbol{J}_{\rm eff}$. This yields $|\boldsymbol{B}_{\rm a}|\sim g_{\rm a\gamma}m_{\rm a}a_{0}\,R\,|\boldsymbol{B}_{0}|$, with $\boldsymbol{B}_{0}$ and $a_{0}$ being respectively the amplitudes of static magnetic and axion fields.
[…]
$|\boldsymbol{B}_{\rm a}|$ $\sim 0.3\,{\rm pT}\left(\frac{g_{\rm a\gamma}}{10^{-10}\,{\rm GeV}^{-1}}\right)\left(\frac{\rho_{\rm DM}}{0.3\,{\rm GeV}\,{\rm cm}^{-3}}\right)^{1/2}\left(\frac{R}{R_{\rm E}}\right)\left(\frac{|\boldsymbol{B}_{\rm geo}|}{50\,\mu{\rm T}}\right).$ (1)
@@
Although the amplitude at Eq. (1) is more than seven orders of magnitude smaller than that of the geomagnetic fields, it is still detectable with high-precision commercial magnetometers. Further, at f=0.330f=0.3-30 Hz, the major background noise is a random superposition of transient EM waves (e.g., Barr et al. (2000); Simões et al. (2012)), which can be discriminated from the static axion signal having a sharp spectral line by using a long-term monitoring data of terrestrial magnetic fields.
@@
Figure 1: Atmospheric conductivity profiles as a function of altitude above the Earth’s surface, $r-R_{\rm E}$. While the solid curve is taken from Ref. Kudintseva et al. (2016), dashed, dash-dotted, and dotted curves correspond to the day-side and night-side profiles of Ref. Galuk et al. (2018) and the median profile of Ref. Nickolaenko et al. (2016), respectively. Note that these profiles are provided as tabulated data with different altitude ranges.
Introducing the dimensionless variable $x\equiv m_{\rm a}r$, this is obtained by solving the following equation under the boundary condition that the tangential electric field vanishes at the Earth’s surface and the homogeneous solution approaches an upward-propagating wave near the lower ionosphere ($r\sim R_{\rm E}+100\,\,\rm{km}$), and retaining only the homogeneous part of the solution (see Ref. Taruya et al. (2025) for more details):
@@
Note that Eq. (3) is derived by using the wave equation for the electric field alongside the Maxwell-Faraday equation. Here, the influence of the atmospheric conductivity is incorporated into the complex refractive index $n$ through $n^{2}=1+i\,(\sigma/m_{\rm a})$, where the conductivity $\sigma$ increases monotonically with the radial coordinate $r$, from $10^{-14}$ to $10^{-3}$ S/m between the Earth’s surface and the lower ionosphere, as shown in Fig. 1 (see Ref. Taruya et al. (2025) for details).
Figure 2: Top: Radial mode function $b_{\ell}$ for the lowest six multipoles, evaluated at the Earth’s surface $r=R_{\rm E}$. The solid curves are the results adopting the conductivity profile of Ref. Kudintseva et al. (2016), which are independent of geographic location, and plotted against the axion mass $m_{\rm a}$ (lower) and the induced EM wave frequency $f_{\rm a}$ (upper). For comparison, the dotted curves indicate the mode functions in an idealized setup that neglects the atmospheric conductivity. Middle: Expected amplitude of the axion-induced magnetic field at Eskdalemuir observatory, normalized by the axion-photon coupling strength of $10^{-10}\,$GeV-1. Following Eq. (2), this is obtained by summing the mode functions from the top panel, weighted by the harmonic coefficients of geomagnetic field and vector spherical harmonics, and evaluating at the observatory’s location. Bottom: Impact of different conductivity profiles on the axion-induced magnetic field, shown as fractional differences. Dashed, dash-dotted, and dotted curves correspond to those shown in Fig. 1, relative to the reference model in the middle panel.
Adopting the conductivity profile in Ref. Kudintseva et al. (2016) (black solid curve in Fig. 1), we solve the radial mode equation by treating the radial dependence of $\sigma$ or $n$ as a sum of spatially constant segments. This approach allows us to analytically solve for $b_{\ell}$, while satisfying the boundary conditions at the Earth’s surface and the lower ionosphere Taruya et al. (2025). The top panel of Figure 2 plots the resultant mode functions for the lowest six multipoles at the Earth’s surface, given as a function of $m_{\rm a}$ or the frequency $f_{a}$ of the induced EM signal. There is a prominent peak structure in each multipole, known as the Schumann resonance Schumann (1952a, b) (see also Refs. Jackson (1998); Nickolaenko and Hayakawa (2002); Simões et al. (2012); Nickolaenko and Hayakawa (2013)), which is caused by EM waves trapped between the Earth’s surface and the lower ionosphere, forming a resonant cavity. The peak frequencies are close to $\sqrt{\ell(\ell+1)}/(2\pi\,R_{\rm E})\simeq 7.46\sqrt{\ell(\ell+1)}$ Hz, as predicted in the idealized case of a perfectly conducting Earth-ionosphere cavity (dotted curves), although the actual peak positions depicted as solid curves are systematically lower than these values by 20-30% due to losses in the propagation of EM waves. As a result, the resonant peaks have a finite amplitude, approximately $b_{\ell}\sim 5$ for $\ell=1-6$, unlike the idealized case where the amplitudes exhibit a divergent behavior.
@@
The middle panel of Fig. 2 shows the predicted magnetic field amplitude at the Eskdalemuir observatory, UK, $(55.31^{\circ}\,{\rm N},\,3.21^{\circ}\,{\rm W})$, where long-term monitoring data up to 100100\,Hz is available (see below). Here, we assume that the axions constitute the DM having the local density of $\rho_{\rm DM}=0.3\,{\rm GeV}\,{\rm cm}^{-3}$, with the coupling parameter of $g_{a\gamma}=10^{-10}$ GeV-1. The induced magnetic field is predicted to exhibit a peak at $m_{\rm a}\sim 3\times 10^{-14}$ eV, corresponding to the peak frequency $f_{\rm a}\sim 7.8$ Hz for the radial mode function of $\ell=1$. Additionally, a low-frequency plateau of $|\boldsymbol{B}_{\rm a}|\sim\mathcal{O}(10^{-1})$ pT is observed, consistent with Ref. Arza et al. (2022). On the other hand, at $m_{\rm a}\gtrsim 10^{-13}$ eV, its amplitude is sharply suppressed, reflecting the dipole nature of the geomagnetic fields, where the higher multipoles are sufficiently small.
The spectral features seen in the middle panel of Fig. 2 appear typical regardless of the geographic location, though the overall amplitude varies mostly with latitude, increasing near the equator, which reflects a toroidal field configuration Taruya et al. (2025). While these predictions rely on the specific atmospheric conductivity model of Ref.Kudintseva et al. (2016), we find that the choice of conductivity profile only has a mild effect. As shown in the bottom panel of Fig.2, even though conductivity models of Refs. Nickolaenko et al. (2016); Galuk et al. (2018) differ by nearly an order of magnitude around the lower ionosphere (see Fig. 1), the actual impact on the magnetic field at the Earth’s surface remains small, reaching at most $\sim 20$% in a regime where the magnetic field is already suppressed.
@@
The basic data analysis is summarized as follows. First, a continuous dataset spanning more than one month is extracted from the full data, which is then divided into segments of length $T_{\rm seg}=8\,{\rm hrs}$. Second, each data segment is Fourier-transformed, and its PSD is computed. The frequency bin size of each segment is $\Delta f_{\rm a}\simeq 1/T_{\rm seg}\simeq 3.5\times 10^{-5}$ Hz, roughly corresponds to the bandwidth of the axion-induced signal at $f_{\rm a}=35$ Hz, above which the signal decays rapidly as shown in the middle panel of Fig. 2. Axion-induced signals at $f_{\rm a}\lesssim 35\,\,\rm{Hz}$ are well within a single frequency bin. Considering the coherence time, the segment size of 8 hrs may not be an optimal choice but is chosen so as to avoid unknown-cause short-duration transient noise with huge amplitude of $\sim 10^{8}\,\rm{pT}$ that happens irregularly. Due to the transient noise, the fraction of the data segments with good quality is roughly $\sim 10$% of the total number of the data segments, $N_{\rm seg}=9909$.
@@
Constraining axion-photon coupling. Given the PSD of the magnetic fields, we stack them with the weight of the inverse of noise variance computed from the spectrum at surrounding frequencies. Then we subtract the smooth component, obtained by applying a quadratic Butterworth filter with a sampling frequency of $0.05\,\rm{Hz}$. This allows the axion-induced line signals to add constructively, while reducing other components (magnetometer noise variance and the Schumann resonance spectrum). Dividing by the sum of the inverse-of-noise-variance weight for data segments, we obtain the weighted average differential PSD, $\hat{s}(f_{\rm a})$. The procedure above still leads to numerous lines at integer-multiple frequencies, the exact origin of which is unknown. Since these are likely artificial, we remove the data surrounding these frequencies ($\pm 1000$ bins). The resultant PSD is shown in Fig. 3.
@@
Figure 3: Weighted-average differential PSD of the magnetic fields, stacked over all eight-hour data segments with noise-based weights and subtracted the smooth spectrum, $\hat{s}(f_{\rm a})$ in blue. The standard deviation is depicted in orange. The axion signal candidates (SNR>3>3) are marked by diamonds in cyan. The inset highlights the signal candidate with the highest SNR $(13.342$).
@@
The search for the axion-induced signal consists of evaluating the signal-to-noise ratio (SNR) for each frequency bin and selecting the candidates that exceed a certain threshold. The standard deviation of the weighted-average differential PSD is calculated from the surrounding $\pm 200$ bins, excluding the three bins around $f=f_{\rm a}$, as shown by the orange curve in Fig. 3. We regard this as noise amplitude to estimate the SNR for the axion signal at $f_{\rm a}$, and evaluate the SNR from $0.01$ Hz, below which the number of samples is insufficient, to $44$ Hz, above which the estimated spectrum becomes less reliable because it is close to the Nyquist frequency.
@@
We then identify signal candidates by requiring (i) ${\rm SNR}>2$ in all-year data, (ii) $\rm SNR$} in each year’s data larger than the SNR threshold (2 in our case) weighted by the sum of the inverse of noise variances in the year, ensuring persistence over the entire observation period, and (iii) that their frequencies are at least $10^{-3}\,\rm{Hz}$ away from multiples of $0.05\,\rm{Hz}$. The third condition is for excluding likely artificial line noises. We found $342$ candidates (see also Nishizawa et al. (2025) for signal candidates with different SNR thresholds). These candidates appear in a single frequency bin, consistent with the sharpness of an axion signal. There are 1 candidate with significant SNR (13.342) and 31 candidates with SNR>5>5 for the all-year data. In Fig. 3, representative candidates of SNR>3>3 are depicted by cyan-filled diamonds. Lacking enough information to veto them, we retain them as potential axion signal candidates.
@@
We note that below $1$ Hz, Ref.  Friel et al. (2024) reported three signal candidates with modest significance using SuperMAG 1-second sampling data. Although our analysis identified five candidates with SNR>2>2, none of them matched those reported in Ref. Friel et al. (2024). Given that our measurement sensitivity is comparable to theirs [^1] To be precise, our sensitivity is slightly worse than theirs for the candidate at 0.26300.2630Hz identified by Ref.  Friel et al. (2024)., the signal candidates identified below $1$ Hz in both analyzes can be ruled out.
@@
Apart from the signal candidates mentioned above, no axion-like signal was found. We can therefore place upper limits on the axion coupling strength in the frequency range of 0.0143.970.01-43.97 Hz. The 95% C. L. upper limit on the axion-photon coupling strength is obtained from
@@
The $\hat{s}_{\rm obs}$ is the observed value of the weighted-average differential PSD, which can be converted to the axion coupling strength by equating it with $\langle\hat{s}_{\rm a}(m_{\rm a})\rangle\equiv 2T_{\rm seg}|\boldsymbol{B}_{\rm a}(m_{\rm a})|^{2}$ (see Ref. Nishizawa et al. (2025) for the derivation) and using the theoretical curve in the middle panel of Fig. 2. The $p[\hat{s}(m_{\rm a})|g_{\rm a\gamma}]$ is the probability density distribution of the weighted-average differential PSD in the presence of axions, constructed from data $\hat{s}$ of the surrounding $\pm 200$ bins (excluding the three bins around $f=f_{\rm a}$) and shifted by the axion signal $\hat{s}_{\rm a}$ at $f=f_{\rm a}$.
@@
Figure 4 shows the upper limit on the axion coupling we obtained (in blue), along with other observation bounds. Tightest limit is obtained around an axion mass of $3\times 10^{-14}\,\rm{eV}$, which is improved by about two orders of magnitude compared to the CAST result  Anastassopoulos et al. (2017); Altenmüller et al. (2024), and exceeds the limit from astrophysical X-ray observations by Chandra Reynés et al. (2021); Reynolds et al. (2020); Dessert et al. (2022); Ning et al. (2025) and NuSTAR Ning and Safdi (2025). In the lower mass range, our constraint is weakened because of fixing the data segment size to 8 hrs and using broader frequency resolution, irrespective of the coherence time of axions. In the higher mass range above $10^{-14}\,\rm{eV}$, the constraint is also weakened due to the suppression of the magnetic field’s response to axions.
@@
Figure 4: Constraint (95% C. L.) on the axion-photon coupling $g_{\rm a\gamma}$ from the long-term monitoring data of magnetic fields at Eskdalemuir observatory (this work, blue) except for the notched frequency bands. Other excluded regions are from SuperMAG Arza et al. (2022); Friel et al. (2024), SNIPE Sulai et al. (2023), Planck and unWISE blue galaxy sample Goldstein et al. (2024), CAST Anastassopoulos et al. (2017); Altenmüller et al. (2024), Chandra Reynés et al. (2021); Reynolds et al. (2020), and NuSTAR Ning and Safdi (2025). The data are taken from Axi . The constraints by direct means are filled and those by astrophysical model-dependent means are not filled.
@@
Conclusions and outlook. In this Letter, we properly accounted for atmospheric conductivity to predict terrestrial EM waves induced by coherently oscillating axions, enabling a search for axion DM signals in public magnetic field data. We set the most stringent direct upper bound on the axion-photon coupling parameter at $10^{-15}\,{\rm eV}\lesssim m_{\rm a}\lesssim 10^{-13}\,{\rm eV}$, and identify potential signal candidates with significant SNR for further study. Follow-up ground-based experiments such as DANCE Obata et al. (2018) and those employing twisted anyon cavities Bourhill et al. (2023), as well as X-ray observations by Athena Sisk-Reynés et al. (2023) will be crucial to confirm or rule out these candidates. While we focus on axions as a representative ultralight DM, our methodology applies to other DM coupled to EM fields, such as dark photon DM Nelson and Scholtz (2011); Fabbrichesi et al. (2020), whose kinetic mixing can be constrained using the same dataset Fedderke et al. (2021a, b); Nomura et al. (2025).